OpenStudy (anonymous):

There are 20 people who work together. Four of these people are selected to go to the same conference together. How many such selections are possible? Select one: a. 80 b. 4,845 c. 116,280 d. None of the above

OpenStudy (anonymous):

This is way i would answer this... 20/4*20/4*20/4*20/4=160,000/4= 40,000 possible choices I am not sure if this is correct but i hope this helped in some way! @phi please help out this question :)

OpenStudy (anonymous):



OpenStudy (anonymous):

@Fransisco_w I checked your answer its correct

OpenStudy (anonymous):

so i'd say none of the above

OpenStudy (anonymous):

It would None of the above then^^^^

OpenStudy (anonymous):

-w-

OpenStudy (anonymous):

Thank You Guys So Much :)

OpenStudy (anonymous):

No problem, Have a nice day!

OpenStudy (phi):

It would 20 choices for the first person, 19 for the 2nd, 18 for the 3rd, and 17 for the last 20*19*18*17